OpenStudy (anonymous):

Chef Imelda can do something unique. Using a secret process, she can bake a nearly perfectly spherical pie consisting of a chicken filling inside a thick crust. The radius of the whole pie is 10 cm, and the radius of the filling is 5 cm. What is the volume of the crust alone? Use 1100-11-09-00-00_v2_files/i0050000.jpg = 3.14

OpenStudy (anonymous):

Volume of the whole pie: \[V=\frac{4}{3}\pi r^{3}\] \[V=\frac{4}{3}\pi 10^{3}\] \[V=\frac{4000}{3} \pi\] Volume of the filling: \[V=\frac{4}{3}\pi r^{3}\] \[V=\frac{4}{3}\pi 5^{3}\] \[V=\frac{500}{3} \pi\] Volume of crust only:\[V_c=V_p-V_f=\frac{3500}{3} \pi\]